OpenStudy (anonymous):

How does 9^((log 5)/(log 3)) equal 25?!?

OpenStudy (anonymous):

Can you please explain? I would be soooooo thankful if somebody could help me

OpenStudy (anonymous):

change of base formula

OpenStudy (anonymous):

\[3^x=5\iff x=\frac{\log(5)}{\log(3)}\]

OpenStudy (anonymous):

i know that log5/log is change of base but how do you raise 9 to a log

OpenStudy (anonymous):

By the way, thank you sooo mcuch for wanting to help me :)

OpenStudy (anonymous):

so if \(3^x=5\) then \(9^x=25\)

OpenStudy (anonymous):

that is clear right?

OpenStudy (anonymous):

But why 3^x where do you get that from im sooo sorry im really confused i just want to understand hw to do it

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

is it clear that \[b^x=A\iff x=\frac{\log(A)}{\log(b)}\]?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

ok so if we put \(b=3, A=5\) we have \[3^x=5\iff x=\frac{\log(5)}{\log(3)}\] right?

OpenStudy (anonymous):

yes, im following :D

OpenStudy (anonymous):

so if \(3^x=5\) then \(9^x=25\) yes?

OpenStudy (anonymous):

YES cause 9 is 3 squared right?!

OpenStudy (anonymous):

as \(3=3^2\) so \(9^x=3^{2x}=5^2=25\)

OpenStudy (anonymous):

THANK YOU SO MUCH!! ahhh you helped me so much

OpenStudy (anonymous):

seriosuly :D

OpenStudy (anonymous):

or to make it even more obvious \[3^x=5\] so \[9^x=(3^2)^x=(3^x)^2=5^2=25\]

OpenStudy (anonymous):

Thank you :) you are awesome

OpenStudy (anonymous):

and so if \[3^{\frac{\log(5)}{\log(3)}}=5\] then \[9^{\frac{\log(5)}{\log(3)}}=5^2=25\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

btw there might be some other way to see this with logs there is usually more than one way to skin a cat

OpenStudy (anonymous):

got it PERFECTLY yes, logs have many ays of a pproaching problems but this was more than enough